Now, assume that Best Buy’s current price is in equilibrium and is a constant growth stock. What is Best Buy’s expected constant growth rate based on your CAPM return? Use price and dividend data from #6.

Stock Best Buy Activision
Current Price 68.40 55.40
Estimated Price from 1 year now 75.89 57.52
Annual Expected Dividend (D1) 2.00 0.37

D2=2.20 P1=75.89 Growth rate (r)=12.90%
Answer #1

Best Buy :

Estimated price from 1 year now = Current price(1+g)

    75.89 = 68.40 (1+g)

    75.89/68.40 = (1+g)

    (1+g) = 1.1095

        g = 1.1095 -1

          = .1095 or 10.95%

constant growth = 10.95%